# Message-Passing Monte Carlo: Generating low-discrepancy point sets via Graph Neural Networks

### v1

@@ Abstract
Discrepancy is a well-known measure for the irregularity of the distribution of a point set. Point sets with small discrepancy are called low-discrepancy and are known to efficiently fill the space in a uniform manner. Low-discrepancy points play a central role in many problems in science and engineering, including numerical integration, computer vision, machine perception, computer graphics, machine learning, and simulation. In this work, we present the first machine learning approach to generate a new class of low-discrepancy point sets named Message-Passing Monte Carlo (MPMC) points. Motivated by the geometric nature of generating low-discrepancy point sets, we leverage tools from Geometric Deep Learning and base our model on Graph Neural Networks. We further provide an extension of our framework to higher dimensions, which flexibly allows the generation of custom-made points that emphasize the uniformity in specific dimensions that are primarily important for the particular problem at hand. Finally, we demonstrate that our proposed model achieves state-of-the-art performance superior to previous methods by a significant margin. In fact, MPMC points are empirically shown to be either optimal or near-optimal with respect to the discrepancy for every dimension and the number of points for which the optimal discrepancy can be determined.
Machine Learning, ICML
1 Introduction
Monte Carlo (MC) methods have been commonly used and are a popular choice for approximating and simulating complex real-world systems. Known for their reliance on repeated random sampling, MC methods function well in problems involving optimization, numerical integration, and financial mathematics (particularly derivative pricing and risk management) via computer simulation. However, their convergence rate of $\mathcal{O}(N^{-1/2})$ in the number of samples $N$ means that achieving high precision with MC requires an impractically large number of samples for complex problems. To address this drawback, it is common to employ variance reduction techniques such as importance sampling, stratified sampling, or control of variates to obtain the same degree of accuracy with fewer samples (for details, see (Glasserman, 2004), (Lemieux, 2009) and references therein).
A particularly successful approach for convergence is called quasi-Monte Carlo (QMC). QMC methods employ a deterministic point set, which replaces the purely random sampling with a sample whose points span the hypercube $[0,1]^{d}$ in a manner that is more uniform than what can be achieved with MC sampling. The fact that these point sets are constructed over $[0,1]^{d}$ is not overly restrictive as most, if not all, sampling algorithms used within the MC method take as input (pseudo)random numbers in $[0,1]$. The uniformity of these deterministic point sets (or indeed, any point set) can be captured by one of several of measures of irregularity of distribution, referred to by the umbrella term discrepancy measures (Drmota & Tichy, 1997). The more uniformly distributed the points are, the lower the discrepancy is; point sets possessing a small enough discrepancy value are called low-discrepancy. In the classical setting, the star-discrepancy, widely regarded as the most important uniformity measure, of an $N-$element point set $\{{\bf X}_{i}\}_{i=1}^{N}$ contained in $[0,1]^{d}$ represents the largest absolute difference between the volume of a test box and the proportion of points of $\{{\bf X}_{i}\}_{i=1}^{N}$ that fall inside the test box,
$D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right):=$ (1)
$\sup_{\mathbf{x}\in[0,1]^{d}}\left|\frac{\#\left(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,\mathbf{x})\right)}{N}-\mu([0,\mathbf{x}))\right|$
where #($\#(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,{\bf x}))$ counts how many points of $\{{\bf X}_{i}\}_{i=1}^{N}$ fall inside the box $[0,{\bf x})=\prod_{i=1}^{d}[0,x_{i})$ for ${\bf x}=(x_{1},\ldots,x_{d})\in[0,1]^{d}$, and $\mu(\cdot)$ denotes the usual Lebesgue measure. Despite the framing in (1), computation of the star-discrepancy is known to be a discrete problem; see (Niederreiter, 1972). The discrepancy is closely related to worst-case integration error of a particular class of functions with the most well-known result being the Koksma-Hlawka inequality; see (Kuipers & Niederreiter, 1974; Hlawka, 1984). Explicitly, given a point set $\{{\bf X}_{i}\}_{i=1}^{N}$ contained in $[0,1]^{d}$, we have
@@ 1 Introduction
where $V(f)$ denotes the variation of the function $f$ in the sense of Hardy and Krause. This result illustrates that points with small discrepancy induce approximations with small errors. Thus in summary, it is of general interest to find $N-$point configurations with smallest discrepancy; see (Cauwet et al., 2020; Galanti & Jung, 1997; Paulin et al., 2022; Mishra & Rusch, 2021; Longo et al., 2021) for examples of QMC implementation.
@@
Given this context, our main goal is to present a machine learning framework that generates point sets with minimal discrepancy. Based on the geometric nature of this problem, we suggest to leverage graph-learning models from Geometric Deep Learning (Bronstein et al., 2021) to achieve this. More concretely, we construct a computational graph based on nearest neighbors of the initial input points and process the encoded input points with a deep message-passing neural network, which is trained to minimize a closed-form solution of a specific discrepancy measure of its decoded and clamped outputs. We term the resulting low-discrepancy points Message-Passing Monte Carlo (MPMC) points. While previous methods are either far from obtaining optimal discrepancy values or intractable to compute (and thus are only available in $d=2,3$ and very small number of points $N\leq 20$), MPMC reaches near-optimal discrepancy in near-real time for potentially thousands of points. This advancement represents a significant step forward in the development of highly efficient sampling methods, which are crucial for many applications in science and engineering. Concrete examples include problems in financial mathematics (L’Ecuyer, 2009), path and motion planning in robotics (Branicky et al., 2001), and enhanced training of $3$-D computer-vision models like Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021).
@@ 2 Background and previous work
Our general goal in this paper is to provide a method for generating point sets with small discrepancy. In the following, we use the term sequence to refer to an infinite series of points, and point set for a finite one. Both these objects are closely related as many results on sequences in dimension $d$ correspond to those on sets in dimension $d+1$; see (Roth, 1954).
A sequence of points $\{{\bf X}_{i}\}_{i=1}^{\infty}$ contained in $[0,1]^{d}$ is called a low-discrepancy sequence if the star-discrepancy of the first $N$ points satisfies $D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right)=\mathcal{O}((\log N)^{d}/N)$. A finite point set $\{{\bf X}_{i}\}_{i=1}^{N}$ is said to be of low discrepancy if its corresponding star-discrepancy $D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right)$ is “small” enough, which in practice means that a bound of the form $c(\log N)^{d-1}/N$ can be established, for a given constant $c$ independent of $N$ (but possibly dependent on $d$). Moving forward, for comparison purposes, we will regularly truncate various known infinite low-discrepancy sequences resulting in a finite point set, which inherits the low-discrepancy property from the underlying infinite sequence.
@@
Figure 1 illustrates two examples of low-discrepancy point sets. On the left-hand side we have a Korobov lattice (Korobov, 1963), which is an example of a lattice rule (Haber, 1970; Sloan & Joe, 1994; Nuyens, 2014; Dick et al., 2022), and on the right-hand side, we see the first 64 points of the two-dimensional Sobol’ sequence (Sobol’, 1967). This construction leverages a widely used building block for many low-discrepancy sequences known as the van der Corput sequence in base $b$ (van der Corput, 1935). It is also an example of what are modernly known as digital $(t,s)$-sequences—which also include the Faure sequences (Faure, 1982)—that were first laid out in (Niederreiter, 1987), with a comprehensive overview provided in the subsequent monograph (Niederreiter, 1992). Halton sequences (Halton, 1960) are another widely used type of low-discrepancy sequences that concatenate $d$ van der Corput sequences in different bases, usually taken as the first $d$ prime numbers.
More recently, there have been successful attempts to construct low-discrepancy point sets using more sophisticated means motivated by the lack of constructions adapted to specific $N$ and $d$. In (Doerr & De Rainville, 2013), new low-discrepancy point sets were suggested via the optimization of permutations applied to a Halton sequence. As a consequence, several open problems were solved regarding sets with small discrepancy from (Novak & Woźniakowski, 2010). Further, in (Doerr et al., 2005) an algorithm for constructing a low-discrepancy set via a derandomized version of Hoeffding’s inequality was provided which was shown to improve the previous best known upper bounds for the star-discrepancy. As some of the most recent work in this direction, a method called subset selection was presented in (Clément et al., 2022, 2024) to choose from an $N-$element point set (in practise, usually the first $N$ points of the Sobol’ sequence) the $k<N$ points which yield the smallest discrepancy using a swap-based heuristic. Furthermore, a method to generate optimal star-discrepancy point sets for fixed $N$ and $d$ based on a non-linear programming approach was suggested in (Clément et al., 2023). However, this formulation of the problem presented huge computational burdens allowing optimal sets only to be found up to $20$ points in dimension two and $8$ points in dimension three.
@@ 3 Method
In this work, we propose to leverage Graph Neural Networks (GNNs) (Sperduti, 1994; Goller & Kuchler, 1996; Sperduti & Starita, 1997; Frasconi et al., 1998; Gori et al., 2005; Scarselli et al., 2008; Bruna et al., 2014; Defferrard et al., 2016; Kipf & Welling, 2017; Monti et al., 2017) based on the message-passing framework to effectively learn such transformations. GNNs are a popular class of model architectures for learning on relational data, and have successfully been applied on a variety of different tasks, e.g., in computer science (Monti et al., 2017; Derrow-Pinion et al., 2021; Ying et al., 2018), and the natural sciences (Gilmer et al., 2017; Gaudelet et al., 2021; Shlomi et al., 2020) (see (Zhou et al., 2019; Bronstein et al., 2021) for additional applications). In particular, GNNs have successfully been used in the context of learning on point clouds, or generally learning on sets. This motivates the choice of GNNs in our setup, where specific transformations of geometric sets (i.e., set of input points in $[0,1]^{d}$) have to be learned.
@@ 3.1 Training set
Our approach can be classified as an unsupervised learning setup, where, in contrast to supervised learning, only input data is required without any labels. While it is intuitive to generate the set of input points randomly, we suggest several different approaches for constructing input data:
  1. 1.
    Uniform random sampled set of input points: ${\bf X}_{i}\sim\mathcal{U}([0,1]^{d})$, for all points $i=1,\dots,N$.
  2. 2.
    Base set of input points from available low-discrepancy point sets, such as Sobol’, Halton, or a lattice rule.
  3. 3.
    Base set of input points from randomly perturbed low-discrepancy points, i.e.,
@@
    where ${\bf Y}_{i}$ is generated by a known low-discrepancy sequence and $\xi$ is uniform randomly sampled from $[0,b]^{d}$, with $0<b\leq 1$, for all points $i=1,\dots,N$.
3.2 Architecture
Let $\mathcal{G}=(\mathcal{V},\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V})$ be an undirected, connected graph with $|\mathcal{V}|=N$ nodes and $|\mathcal{E}|=e$ edges (unordered pairs of nodes $\{i,j\}$ denoted $i\sim j$). We further denote the $1$-neighborhood of a node $i\in\mathcal{V}$ as $\mathcal{N}_{i}=\{j\in\mathcal{V}:i\sim j\}$. In addition, each node $i\in\mathcal{V}$ is equipped with an $m$-dimensional feature vector ${\bf X}_{i}\in{\mathbb{R}}^{m}$. The main building block of our model consist of GNN layers based on the message-passing framework. This family of parametric functions is defined through local updates of hidden node representations. More concretely, we iteratively update node features as,
${\bf X}_{i}^{l}=\phi^{l}\left({\bf X}_{i}^{l-1},\bigoplus_{j\in\mathcal{N}_{i}}\psi^{l}({\bf X}_{i}^{l-1},{\bf X}_{j}^{l-1})\right),~{}\forall l=1,\dots,L,$ (4)
where $\bigoplus$ denotes a permutation-invariant operation, such as $\mathsf{SUM}$, $\mathsf{MEAN}$, or $\mathsf{MAX}$, and ${\bf X}_{i}^{l}\in{\mathbb{R}}^{m_{l}}$ for all nodes $i$. Moreover, we parameterize $\phi^{l},\psi^{l}$ as ReLU-multilayer perceptrons (MLPs), i.e., MLPs using the element-wise $\text{ReLU}(x)=\max(0,x)$ activation function in-between layers. We further encode the initial node features by a transformation that maps the initial points in dimension $d$ to our initial hidden node feature dimension $m_{0}$. Moreover, we decode the output of the final GNN layer by a transformation that maps the hidden node features of dimension $m_{L}$ back to the physical dimension $d$. Both the encoder as well as the decoder are parameterized as affine transformations. Finally, we smoothly clamp the decoded outputs back into $[0,1]^{d}$ by using the element-wise sigmoidal activation function,
$\text{sigmoid}(x)=\frac{1}{1+e^{-x}}.$
Note that this step is crucial, as otherwise the training objectives we introduce in the subsequent sections are ill-defined. Moreover, the clamping step has to be differentiable in order to be used within a gradient-based learning framework.
Another important part of the GNN architecture is the construction of the underlying computational graph $\mathcal{G}$, i.e., defining the local structure $\mathcal{N}_{i}$ for all nodes $i$ in (4). It is worth noting that in many GNN applications (e.g., network science, or life sciences) the computational graph is already given a-priori, either explicitly or implicitly. In contrast to that, our problem setup considers the construction of the underlying computational graph as an additional design choice. While there are many suitable choices, often balancing a global vs local connectivity structure, we suggest to construct the underlying graph $\mathcal{G}$ based on nearest neighbors, i.e., for a fixed radius $0<r\leq 1$,
@@ 3.2 Architecture
We choose this inherently local structure to guide the GNN training towards transforming input points into low-discrepancy points by mainly considering the positions of other near-by points (in the corresponding Euclidean space of the input point set). A schematic of the full model can be seen in Fig. 3.

### v2

# pnasresearcharticle

@@ Abstract
Discrepancy is a well-known measure for the irregularity of the distribution of a point set. Point sets with small discrepancy are called low-discrepancy and are known to efficiently fill the space in a uniform manner. Low-discrepancy points play a central role in many problems in science and engineering, including numerical integration, computer vision, machine perception, computer graphics, machine learning, and simulation. In this work, we present the first machine learning approach to generate a new class of low-discrepancy point sets named Message-Passing Monte Carlo (MPMC) points. Motivated by the geometric nature of generating low-discrepancy point sets, we leverage tools from Geometric Deep Learning and base our model on Graph Neural Networks. We further provide an extension of our framework to higher dimensions, which flexibly allows the generation of custom-made points that emphasize the uniformity in specific dimensions that are primarily important for the particular problem at hand. Finally, we demonstrate that our proposed model achieves state-of-the-art performance superior to previous methods by a significant margin. In fact, MPMC points are empirically shown to be either optimal or near-optimal with respect to the discrepancy for low dimension and small number of points, i.e., for which the optimal discrepancy can be determined. Code for generating MPMC points can be found at https://github.com/tk-rusch/MPMC
keywords:
Low-discrepancy, Machine Learning, Graph Neural Networks, Geometric Deep Learning, Quasi-Monte Carlo
21]2
@@
Monte Carlo (MC) methods have been commonly used and are a popular choice for approximating and simulating complex real-world systems. Known for their reliance on repeated random sampling, MC methods function well in problems involving optimization, numerical integration, and financial mathematics (particularly derivative pricing and risk management) via computer simulation. However, their convergence rate of $\mathcal{O}(N^{-1/2})$ in the number of samples $N$ means that achieving high precision with MC requires an impractically large number of samples for complex problems. To address this drawback, it is common to employ variance reduction techniques such as importance sampling, stratified sampling, or control variates to obtain the same degree of accuracy with fewer samples (for details, see (1), (2) and references therein).
A particularly successful approach for convergence is called quasi-Monte Carlo (QMC). QMC methods employ a deterministic point set, which replaces the purely random sampling with a sample whose points span the hypercube $[0,1]^{d}$ in a manner that is more uniform than what can be achieved with MC sampling. The fact that these point sets are constructed over $[0,1]^{d}$ is not overly restrictive as most, if not all, sampling algorithms used within the MC method take as input (pseudo)random numbers in $[0,1]$. The uniformity of these deterministic point sets (or indeed, any point set) can be captured by one of several of measures of irregularity of distribution, referred to by the umbrella term discrepancy measures (3). The more uniformly distributed the points are, the lower the discrepancy is; point sets possessing a small enough discrepancy value are called low-discrepancy. In the classical setting, the star-discrepancy, widely regarded as the most important uniformity measure, of an $N-$element point set $\{{\bf X}_{i}\}_{i=1}^{N}$ contained in $[0,1]^{d}$ represents the largest absolute difference between the volume of a test box and the proportion of points of $\{{\bf X}_{i}\}_{i=1}^{N}$ that fall inside the test box,
$D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right):=\sup_{\mathbf{x}\in[0,1]^{d}}\left|\frac{\#\left(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,\mathbf{x})\right)}{N}-\mu([0,\mathbf{x}))\right|$ (1)
where #($\#(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,{\bf x}))$ counts how many points of $\{{\bf X}_{i}\}_{i=1}^{N}$ fall inside the box $[0,{\bf x})=\prod_{i=1}^{d}[0,x_{i})$ for ${\bf x}=(x_{1},\ldots,x_{d})\in[0,1]^{d}$, and $\mu(\cdot)$ denotes the usual Lebesgue measure. The discrepancy is closely related to worst-case integration error of a particular class of functions with the most well-known result being the Koksma-Hlawka inequality; see (4, 5). Explicitly, given a point set $\{{\bf X}_{i}\}_{i=1}^{N}$ contained in $[0,1]^{d}$, we have
[…]
@@
where $V(f)$ denotes the variation of the function $f$ in the sense of Hardy and Krause. This result illustrates that points with small discrepancy induce approximations with small errors. Thus in summary, it is of general interest to find $N-$point configurations with smallest discrepancy; see (6, 7, 8, 9, 10) for examples of QMC implementation.
@@
Given this context, our main goal is to present a machine learning framework that generates point sets with minimal discrepancy. Based on the geometric nature of this problem, we suggest to leverage graph-learning models from Geometric Deep Learning (11) to achieve this. More concretely, we construct a computational graph based on nearest neighbors of the initial input points and process the encoded input points with a deep message-passing neural network, which is trained to minimize a closed-form solution of a specific discrepancy measure of its decoded and clamped outputs. We term the resulting low-discrepancy points Message-Passing Monte Carlo (MPMC) points. Previous methods either fail to achieve optimal discrepancy values or are computationally intractable, being limited to small dimensions ($d\leq 3$) and very small numbers of points ($N\leq 21$), which can still require weeks of computation. In contrast, MPMC can be trained in a few minutes to achieve near-optimal discrepancy for these cases. Moreover, we show that MPMC is not limited to small number of points in small dimensions but can efficiently generate $N>1000$ points in tens of dimensions. This advancement represents a significant step forward in the development of highly efficient sampling methods, which are crucial for many applications in science and engineering. Concrete examples include problems in financial mathematics (12), path and motion planning in robotics (13), and enhanced training of neural scene rendering methods like Neural Radiance Fields (NeRFs) (14).
@@ 1 Background and previous work
Our general goal in this paper is to provide a method for generating point sets with small discrepancy. In the following, we use the term sequence to refer to an infinite series of points, and point set for a finite one.
A sequence of points $\{{\bf X}_{i}\}_{i=1}^{\infty}$ contained in $[0,1]^{d}$ is called a low-discrepancy sequence if the star-discrepancy of the first $N$ points satisfies $D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right)=\mathcal{O}((\log N)^{d}/N)$. A finite point set $\{{\bf X}_{i}\}_{i=1}^{N}$ is said to be of low discrepancy if its corresponding star-discrepancy $D^{*}\left(\{{\bf X}_{i}\}_{i=1}^{N}\right)$ is “small” enough, which in practice means that a bound of the form $c(\log N)^{d-1}/N$ can be established, for a given constant $c$ independent of $N$ (but possibly dependent on $d$). Moving forward, for comparison purposes, we will truncate various known infinite low-discrepancy sequences resulting in a finite point set, which inherits the low-discrepancy property from the underlying infinite sequence.
@@
Figure 1 illustrates two examples of low-discrepancy point sets. On the left-hand side we have a Korobov lattice (15), which is an example of a lattice rule (16, 17, 18, 19), and on the right-hand side, we see the first 64 points of the two-dimensional Sobol’ sequence (20). This construction leverages a widely used building block for many low-discrepancy sequences known as the van der Corput sequence in base $b$ (21). It is also an example of what are modernly known as digital $(t,s)$-sequences—which also include the Faure sequences (22)—that were first laid out in (23), with a comprehensive overview provided in the subsequent monograph (24). Halton sequences (25) are another widely used type of low-discrepancy sequences that concatenate $d$ van der Corput sequences in different bases, usually taken as the first $d$ prime numbers.
More recently, there have been successful attempts to construct low-discrepancy point sets using more sophisticated means motivated by the lack of constructions adapted to specific $N$ and $d$. In (26), new low-discrepancy point sets were suggested via the optimization of permutations applied to a Halton sequence. More recently, a method called subset selection is formulated to choose from an $N-$element point set, the $k<N$ points which yield the smallest discrepancy. An exact selection algorithm was presented in (27), while a swap-based heuristic approach was used in (28). Furthermore, a method to generate optimal star-discrepancy point sets for fixed $N$ and $d$ based on a non-linear programming approach was suggested in (29). However, this formulation of the problem presented huge computational burdens allowing optimal sets only to be found up to $21$ points in dimension two and $8$ points in dimension three.
@@ 2 Method
In this work, we propose to leverage Graph Neural Networks (GNNs) (30, 31, 32, 33, 34, 35, 36, 37, 38, 39) based on the message-passing framework to effectively learn such transformations. GNNs are a popular class of model architectures for learning on relational data, and have successfully been applied on a variety of different tasks, e.g., in computer science (39, 40, 41), and the natural sciences (42, 43, 44) (see (45, 11) for additional applications). In particular, GNNs have successfully been used in the context of learning on point clouds, or generally learning on sets. This motivates the choice of GNNs in our setup, where specific transformations of geometric sets (i.e., set of input points in $[0,1]^{d}$) have to be learned.
@@ 2.1 Training set
Our approach can be classified as an unsupervised learning setup, where, in contrast to supervised learning, only input data is required without any labels. While it is intuitive to generate the set of input points randomly, we suggest several different approaches for constructing input data. First, using aforementioned uniform random sampled set of input points ${\bf X}_{i}\sim\mathcal{U}([0,1]^{d})$, for all points $i=1,\dots,N$. Second, using set of input points from available low-discrepancy constructions, such as Sobol’, Halton, or a lattice rule. Third, using set of input points from randomly perturbed low-discrepancy points, i.e.,
@@
where ${\bf Y}_{i}$ is generated by a known low-discrepancy set and $\xi$ is uniform randomly sampled from $[0,b]^{d}$, with $0<b\leq 1$, for all points $i=1,\dots,N$.
2.2 Model architecture
We start by constructing an undirected computational graph $\mathcal{G}=(\mathcal{V},\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V})$, where $\mathcal{V}$ denotes the set of unordered nodes corresponding to the input points $\{{\bf X}_{i}\}_{i=1}^{N}$, and $\mathcal{E}$ is the set of pair-wise connections between the nodes. In addition, each node $i\in\mathcal{V}$ is equipped with a node feature set to the coordinates of an input point, i.e., set to ${\bf X}_{i}\in{\mathbb{R}}^{d}$. We further denote the $1$-neighborhood of a node $i\in\mathcal{V}$ as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$. Clearly, the set of all 1-neighborhoods induces the connectivity of the graph, i.e., the set of node-wise connections $\mathcal{E}$. Hence, the only remaining part of the construction of the underlying computational graph $\mathcal{G}$ is to define the local structure of the graph, i.e., defining $\mathcal{N}_{i}$ for all nodes $i$. It is worth noting that in many GNN applications (e.g., network science, or life sciences) the computational graph structure is already given a-priori, either explicitly or implicitly. In contrast to that, our problem setup considers the connectivity of the underlying computational graph as an additional design choice. While there are many suitable choices, often balancing a global vs local connectivity structure, we suggest basing the local connectivity on the nearest neighbors of the node features, i.e., for a fixed radius $0<r\leq\sqrt{d}$,
@@ 2.2 Model architecture
We choose this inherently local structure to guide the GNN training towards transforming input points into low-discrepancy points by mainly considering the positions of other near-by points (in the corresponding Euclidean space of the input point set).
We can now define the main building block of our MPMC model, i.e., GNN layers based on the message-passing framework. Message-passing GNNs are a family of parametric functions defined through local updates of hidden node representations. More concretely, we iteratively update node features as,
${\bf X}_{i}^{l}=\phi^{l}\left({\bf X}_{i}^{l-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum\limits_{j\in\mathcal{N}_{i}}}\psi^{l}({\bf X}_{i}^{l-1},{\bf X}_{j}^{l-1})\right),\vspace{3mm}\text{for all }l=1,\dots,L,$ (5)
with ${\bf X}_{i}^{l}\in{\mathbb{R}}^{m_{l}}$ for all nodes $i$. Moreover, we parameterize $\phi^{l},\psi^{l}$ as ReLU-multilayer perceptrons (MLPs), i.e., MLPs using the element-wise $\text{ReLU}(x)=\max(0,x)$ activation function in-between layers. We further encode the initial node features by an affine transformation ${\bf X}_{i}^{0}={\bf A}_{\text{enc}}{\bf X}_{i}+{\bf b}_{\text{enc}}$ for all $i=1,\dots,N$, with weight matrix ${\bf A}_{\text{enc}}\in{\mathbb{R}}^{m_{0}\times d}$ and bias ${\bf b}_{\text{enc}}\in{\mathbb{R}}^{m_{0}}$. Finally, we decode the output of the final GNN layer by an affine transformation and smoothly clamp the decoded outputs back into $[0,1]^{d}$ by using the element-wise sigmoidal activation function, i.e., $\hat{{\bf X}}_{i}=\sigma({\bf A}_{\text{dec}}{\bf X}_{i}^{L}+{\bf b}_{\text{dec}})$ for all $i=1,\dots,N$, with sigmoidal function $\sigma(x)=1/(1+e^{-x})$, weight matrix ${\bf A}_{\text{dec}}\in{\mathbb{R}}^{d\times m_{L}}$, and bias ${\bf b}_{\text{dec}}\in{\mathbb{R}}^{d}$. Note that clamping is crucial, as otherwise the training objectives we introduce in the subsequent sections are ill-defined. A schematic of the full model can be seen in Fig. 3.